# Rethinking Table Instruction Tuning

### v1

@@ Abstract
Recent advances in table understanding have focused on instruction-tuning large language models (LLMs) for table-related tasks. However, existing research has overlooked the impact of hyperparameter choices and lacks a comprehensive evaluation of the out-of-domain table understanding ability and the general capabilities of these table LLMs. In this paper, we evaluate these abilities in existing table LLMs, and reveal significant declines in both out-of-domain table understanding and general capabilities compared to their base models. Through systematic analysis, we show that hyperparameters, such as learning rate, can significantly influence both table-specific and general capabilities. Contrary to the existing table instruction-tuning works, we demonstrate that smaller learning rates and fewer training instances can enhance table understanding while preserving general capabilities. Based on our findings, we introduce TAMA, a TAble LLM instruction-tuned from LLaMA 3.1 8B Instruct, which achieves performance on par with, or surpassing GPT-3.5 and GPT-4 on table tasks, while maintaining strong out-of-domain generalization and general capabilities. Our findings highlight the potential for reduced data annotation costs and more efficient model development through careful hyperparameter selection.
@@ 1 Introduction
Recent years have witnessed a paradigm shift to data-driven methods for table understanding. Researchers have instruction-tuned various LLMs, particularly the open-source models from LLaMA family (Touvron et al., 2023; Dubey et al., 2024) to improve their ability on handling table-related tasks (Chen et al., 2019; Nan et al., 2022), and pushing the state-of-the-art performance on various table benchmarks (Zhang et al., 2024a, b).
However, existing research has been influenced by the lack of transparency in closed-source LLMs, which often claim to be trained on large-scale datasets without revealing the detailed training process. As a result, open-source efforts have tended to follow these closed-source models by focusing primarily on large-scale datasets (Zhang et al., 2024a), while overlooking the crucial influence of hyperparameter choices. In addition, existing works lack the discussion of how these table LLMs perform on out-of-domain table understanding tasks, and how they compromise their general ability when specializing on table tasks. We argue that out-of-domain table understanding is crucial for table LLMs, as it reflects how well these models generalize to unseen table tasks. In addition, the general capabilities of these models are still important for handling table-related tasks. For instance, instruction following is crucial in real-world applications where end-users may request specific input-output formats (e.g., The user may request the model to return the answer in JSON). Additionally, stronger reasoning capabilities and comprehensive general knowledge can enhance these models’ ability to handle diverse scenarios, such as interpreting user queries and reasoning over complex data. Therefore, having an understanding of these table LLMs’ general capabilities gives us a comprehensive understanding of these models’ limitations in our practical usage.
@@
We then select the latest LLaMA 3.1 8B Instruct model, and proceed to explore how hyperparameter choices influence the model’s performance. Our analysis reveals that learning rate plays a crucial role in shaping the model’s table understanding ability and influencing the model’s general ability. A large learning rate, as seen in the existing table LLMs, compromises the model’s general capabilities and leads to suboptimal table understanding performance. On the other hand, a small learning rate, while effectively preserving the model’s general capabilities, fails to sufficiently improve its table understanding ability. In addition, we find that it is possible to achieve strong table understanding ability with a much smaller amount of training data – for instance, 2,600 in Section 4. Our training size is significantly smaller compared to the two million instances used by TableLLaMA (Zhang et al., 2024a), and ten times smaller than that of TableBenchLLM (Wu et al., 2024), highlighting the potential to reduce annotation costs in future model development. We also explore the effects of epoch numbers and the task synergy, and discuss our findings in Section 3.
@@
In summary, our contributions are three folds:
@@
  • We reveal the impacts of the often-ignored hyperparameter selection such as the learning rate, number of training instances, etc. We find that the commonly-adopted learning rate can be too large, and may lead to suboptimal table understanding performance and compromises the model’s general capabilities. In addition, we can achieve strong table understanding ability with a much smaller amount of training data compared to the existing works.
  • Based on our findings, with careful hyperparameter selection, we instruction-tune LLaMA 3.1 8B Instruct model with 2,600 table instruction data. As an 8B size model, our resulting model, TAMA achieves performance on par with, or even exceeding GPT-3.5 in table understanding tasks, and in some cases surpasses GPT-4, while retaining the general capabilities of its base model. Moreover, TAMA exhibits strong out-of-domain table understanding and general capabilities (Figure 1).
@@
TableLLM (2024b)
CodeLLaMA 7B
& 13B Instruct
2e-5 6 309K R + S
TableBenchLLM
LLaMA 3.1-8B
& others
2e-5 3 20K S
Table 1: Information of existing table instruction tuned models. For “Data Source”, “S” and “R” represent synthesized data and real data, respectively. $\dagger$: a variant based on the LLaMA 2 7B model.
@@ 2.2 Findings
Existing Table LLMs possess limited out-of-domain table understanding ability. In Table 3, all the existing table LLMs suffer from performance drops on Table-Syn compared to their base models. Though these table LLMs achieve SOTA performance on various benchmarks (Zhang et al., 2024a, b), such a performance decline reveals their limited out-of-domain table understanding capabilities, which aligns with the findings by Zheng et al. (2024).
@@ 3.1 Experimental Setup
We conduct full parameter table instruction tuning using the 8B version of the LLaMA 3.1 Instruct model (Dubey et al., 2024) because of its superior general capabilities, especially its strong instruction following ability. Section C.1 provides detailed reasons for our model choice.
@@ 3.2 Analysis
We find that the learning rate plays a crucial role in determining model performance, as well as how well the model preserves its general capabilities. In general, LLaMA 3.1 8B Instruct achieves the best performance when the learning rate is around 1.0e-6 and 5.0e-7. For instance, on TabFact, LLaMA 3.1 8B Instruct achieves its best performance (73.10) at a learning rate of 1.0e-6 with 1500 examples. Moreover, there is little to no decline in LLaMA 3.1 8B Instruct’s performance on MMLU and IFEval with such learning rates. With a smaller learning rate such as 1.0e-7, though the model’s performance on MMLU and IFEval can be well-preserved, the model’s performance on table tasks such as FEVEROUS is suboptimal under the same setup (66.86 compared to 74.63 at a learning rate of 5.0e-6). In contrast, when the learning rate is too large, such as 1.0e-5, we observe a significant decline in the model’s performance on both MMLU and IFEval, suggesting that a larger learning rate may hurt the model’s general capabilities. We note that all the existing table LLMs use a large learning rate of 2e-5 (Table 1), which explains their compromised out-of-domain table understanding ability and general capabilities compared to their base models in Table 3.
@@
We observe in Figure 2 that on table tasks such as FeTaQA and HiTab, there is a period where the model’s performance boosts up quickly, typically happening when tuning on the first 200 examples. Later, the performance improvement seems marginal. This aligns with the findings from Zhou et al. (2024) that the foundational LLM’s performance can be improved with a limited amount of high-quality data in the instruction tuning stage. We hypothesize that with the first few hundred examples, the model is able to enhance its table reasoning ability quickly. After this point, the model’s performance increase may primarily come from fitting the nuanced patterns in these datasets. Therefore, unlike the existing table LLMs which may involve up to two million training instances as seen in Table 1, we choose to train on 200 instances for each dataset in Section 4.
In addition, we can achieve competitive or even SOTA performance with limited data. On HiTab, with a learning rate of 1.0e-6 and 1,500 examples, we achieve an accuracy score of 66.29, outperforming the previous SOTA performance of 64.71 by TableLLaMA. On FEVEROUS, with 1,500 examples, we achieve a better score of 74.63 compared to 73.77 by TableLLaMA. Though the credit also comes from the LLaMA 3.1 Instruct model, which is much stronger compared to the LLaMA 2 model that TableLLaMA is tuned from, we highlight that TableLLaMA has used two million data points in its table instruction tuning stage, including the entire training set of TabFact, FeTaQA, and HiTab, while here we use around 7% of the entire training data for HiTab. Our analysis demonstrates that with a strong foundational model and a good choice of learning rate, we can achieve competitive performance on table understanding tasks with limited training instances.
We provide analysis on the effects of epochs in Section C.2 and multi-task training in Section C.3. We do not see significant performance gains when we increase the number of epochs, therefore we choose to train our model for two epochs in Section 4. We find that there are synergy effects on these table tasks, therefore we decide to fine-tune our model on a diverse range of tasks in Section 4. We provide further analysis across LLMs in Section C.4, and analysis in terms of LoRA and QLoRA in Section C.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7.
4 TAMA
Based on our findings from Section 3, we start building our general table understanding model, TAMA by instruction tuning the LLaMA 3.1 8B Instruct model.
4.1 Experimental Setup
@@ 4.3 Hindsight Analysis
As shown in Figure 3(a), on the table understanding tasks, the learning rate of 1.0e-6 and 5.0e-7 yield the best overall performance, which coincides with our findings in Section 3. In addition, the model achieves its best aggregated performance around two to three epochs for both learning rate.
@@
For MMLU, both 1.0e-6 and 5.0e-7 maintains their performance, sometimes even slightly better than the original LLaMA 3.1 8B Instruct model. As revealed in Table 8, the performance boost is most pronounced on STEM category. We hypothesize that this is because table-related tasks typically involves data analysis that requires math reasoning, which belongs to the STEM category. Therefore, training on table-related tasks would lead to better STEM performance. This also explains the performance boost for MMLUPro in Table 7.
@@ 5 Related Works
The past decade has witnessed a paradigm shift in approaches to table understanding. Before the advent of LLMs, researchers typically adapt model structures to better interpret table data (Lebret et al., 2016; Liu et al., 2018; Yang et al., 2022). As language models demonstrate promising performance on various tasks (Devlin et al., 2019), researchers gradually shift their attention towards data-driven methods for table understanding. For instance, Yin et al. (2020); Herzig et al. (2020) pre-train BERT (Devlin et al., 2019) or BERT-derived model on large-volume of table data from sources such as Wikipedia to acquire better table representations. Xie et al. (2022) reveal the synergy effects of various structured tasks, including many table tasks, laying foundations to build a generalist model for structured data. In the era of LLMs, as LLMs possess innate table-understanding abilities, researchers also explore prompt engineering techniques to optimize LLMs for table tasks (Chang and Fosler-Lussier, 2023; Deng et al., 2024).
@@
Building on the advances in data-driven methods, researchers have increasingly focused on instruction tuning to enhance LLMs’ table understanding ability. As demonstrated by Touvron et al. (2023); Dubey et al. (2024); Chung et al. (2024); Su et al. (2024), instruction-tuning can improve model performance and generalization to unseen tasks. Meanwhile, models from the open-source LLaMA family (Touvron et al., 2023) demonstrate strong capabilities, leading researchers to instruction-tune these models for better table understanding. For instance, TableLLaMA (Zhang et al., 2024a) is instruction-tuned from a variant of LLaMA 2 model (Touvron et al., 2023), TableLLM (Zhang et al., 2024b) is instruction-tuned from CodeLLaMA, Wu et al. (2024) instruction-tune various foundational models such as LLaMA 3.1 (Dubey et al., 2024), resulting in their TableBenchLLM model. Moreover, Zheng et al. (2024) treat tables as images and instruction-tune Vicuna (Chiang et al., 2023), a vision model that is originally fine-tuned from the LLaMA model, for table understanding. However, as revealed by Zheng et al. (2024); Deng et al. (2024), treating tables as texts rather than images yields better performance. In this paper, we focus on table instruction tuning with tables fed as texts.
@@ 6 Conclusion
In this paper, we reveal the limited out-of-domain table understanding ability and general capabilities of the existing table LLMs. From our analysis, we find that the commonly-adopted hyperparameters in existing table LLMs are suboptimal, and hyperparameter choices in table instruction tuning are crucial in shaping the model’s capabilities. We select hyperparameters from our analysis, and fine-tune our own model, TAMA. Notably, as an 8B model, TAMA demonstrates strong table understanding ability, outperforming GPT-3.5 on most of the table understanding benchmarks, even achieving performance on par or better than GPT-4. Moreover, TAMA preserves strong general capabilities. We hope our findings and our model TAMA can facilitate future research on structured data.
Limitations
Due to the space constraint, we provide further analysis across LLMs in Section C.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section C.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7. However, due to the scope of one study, we cannot exhaust every possible foundational models and every possible dataset.
We want to emphasize the contributions of our work in terms of three aspects. First, we provide a systematic examination of the overspecialization issues in existing table LLMs, which is valuable for the community as it highlights the limitations and potential of the existing models. Second, we investigate how different training configurations influence model behavior. Our findings show that with proper setups, we can maintain a model’s general capabilities while improving its abilities to deal with table tasks, offering practical guidelines for efficient model development. Our systematic exploration has been largely overlooked in the existing literature, and our paper can fill the existing literature gap. Third, our model is a meaningful addition to the open-source LLM community, providing a strong baseline and facilitating future research in domain-specific instruction tuning.

### v2

@@ Abstract
Recent advances in table understanding have focused on instruction-tuning large language models (LLMs) for table-related tasks. However, existing research has overlooked the impact of hyperparameter choices, and also lacks a comprehensive evaluation of the out-of-domain table understanding ability and the general capabilities of these table LLMs. In this paper, we evaluate these abilities in existing table LLMs, and find significant declines in both out-of-domain table understanding and general capabilities as compared to their base models. Through systematic analysis, we show that hyperparameters, such as learning rate, can significantly influence both table-specific and general capabilities. Contrary to the previous table instruction-tuning work, we demonstrate that smaller learning rates and fewer training instances can enhance table understanding while preserving general capabilities. Based on our findings, we introduce TAMA, a TAble LLM instruction-tuned from LLaMA 3.1 8B Instruct, which achieves performance on par with, or surpassing GPT-3.5 and GPT-4 on table tasks, while maintaining strong out-of-domain generalization and general capabilities. Our findings highlight the potential for reduced data annotation costs and more efficient model development through careful hyperparameter selection. We open-source the project and our models.111 Our project page is at https://lit.eecs.umich.edu/TAMA/. We open-source our code at https://github.com/MichiganNLP/TAMA. We open-source our models at https://huggingface.co/MichiganNLP/tama-5e-7 and https://huggingface.co/MichiganNLP/tama-1e-6. We open-source the data used in this paper at https://huggingface.co/datasets/MichiganNLP/TAMA_Instruct.
@@ 1 Introduction
Recent years have witnessed a paradigm shift to data-driven methods for table understanding. Researchers have instruction-tuned various LLMs, particularly the open-source models from the LLaMA family (Touvron et al., 2023; Dubey et al., 2024), to improve their ability on handling table-related tasks (Chen et al., 2019; Nan et al., 2022), and to advance the state-of-the-art performance on various table benchmarks (Zhang et al., 2024a, b).
However, existing research has been influenced by the lack of transparency on closed-source LLMs, which often claim to be trained on large-scale datasets without revealing the detailed training process. As a result, open-source efforts have tended to follow these closed-source models by focusing primarily on large-scale datasets (Zhang et al., 2024a), while overlooking the crucial influence of hyperparameter choices. In addition, existing work lacks a discussion of how these table LLMs perform on out-of-domain table understanding tasks, and how their general capabilities get compromised when specializing on table tasks. We argue that out-of-domain table understanding is crucial for table LLMs, as it reflects how well these models generalize to unseen table tasks. In addition, the general capabilities of these models are still important for handling table-related tasks. For instance, instruction following is crucial in real-world applications where end-users may request specific input-output formats (e.g., The user may request the model to return the answer in JSON). Additionally, stronger reasoning capabilities and comprehensive general knowledge can enhance these models’ ability to handle diverse scenarios, such as interpreting user queries and reasoning over complex data. Therefore, having an understanding of these table LLMs’ general capabilities gives us a comprehensive understanding of these models’ limitations in our practical usage.
@@
We then select the latest LLaMA 3.1 8B Instruct model, and proceed to explore how hyperparameter choices influence the model’s performance. Our analysis reveals that the learning rate plays a crucial role in shaping the model’s table understanding ability and influencing the model’s general ability. A large learning rate, as seen in the existing table LLMs, compromises the model’s general capabilities and leads to suboptimal table understanding performance. On the other hand, a small learning rate, while effectively preserving the model’s general capabilities, fails to sufficiently improve its table understanding ability. In addition, we find that it is possible to achieve strong table understanding ability with a much smaller amount of training data – for instance, 2,600 in Section 4. Our training size is significantly smaller compared to the two million instances used by TableLLaMA (Zhang et al., 2024a), and ten times smaller than that of TableBenchLLM (Wu et al., 2024), highlighting the potential to reduce annotation costs in future model development. We also explore the effects of the number of epochs and the synergy of the task, and discuss our findings in Section 3.
@@
In summary, our contributions are three fold:
@@
  • We reveal the impact of the often-ignored hyperparameter selection such as the learning rate, number of training instances, and so on. We find that the commonly-adopted learning rate can be too large, may lead to suboptimal table understanding performance, and can compromise the model’s general capabilities. In addition, we can achieve strong table understanding ability with a much smaller amount of training data compared to the existing works.
  • Based on our findings, with careful hyperparameter selection, we instruction-tune LLaMA 3.1 8B Instruct model with 2,600 table instruction data. As an 8B size model, our resulting model, TAMA achieves performance on par with, or even exceeding GPT-3.5 in table understanding tasks, and in some cases surpassing GPT-4, while retaining the general capabilities of its base model. Moreover, TAMA exhibits strong out-of-domain table understanding and general capabilities (Figure 1).
@@
TableLLM (2024b) CodeLLaMA 7B & 13B Instruct 2e-5 6 309K R + S
TableBenchLLM (2024) LLaMA 3.1-8B & others 2e-5 3 20K S
Table 1: Existing table instruction tuned models. For “Data Source”, “S” and “R” represent synthesized data and real data, respectively. $\dagger$: a variant based on the LLaMA 2 7B model.
@@ 2.2 Findings
Existing Table LLMs possess limited out-of-domain table understanding ability. In Table 3, all the existing table LLMs suffer from performance drops on Table-Syn compared to their base models. Though these table LLMs achieve SOTA performance on various benchmarks (Zhang et al., 2024a, b), such a performance decline reveals their limited out-of-domain table understanding capabilities, which aligns with the findings by Zheng et al. (2024a).
@@ 3.1 Experimental Setup
We conduct full parameter table instruction tuning using the 8B version of the LLaMA 3.1 Instruct model (Dubey et al., 2024) because of its superior general capabilities, especially its strong instruction following ability. Section C.1 provides detailed reasons for the choice of LLaMA 3.1 and the choice between the base versus the instruction-tuned version of the model. In addition, we expand our analyses across various LLMs and learning setups detailed in Section 3.2.
@@ 3.2 Analysis
We first analyze the effects of the learning rate and the number of examples.
@@
We find that the learning rate plays a crucial role in determining model performance, as well as how well the model preserves its general capabilities. In general, LLaMA 3.1 8B Instruct achieves the best performance when the learning rate is around 1.0e-6 and 5.0e-7. For instance, on TabFact, LLaMA 3.1 8B Instruct achieves its best performance (73.10) at a learning rate of 1.0e-6 with 1500 examples. Moreover, there is little to no decrease in LLaMA 3.1 8B Instruct performance on MMLU and IFEval with such learning rates (Figures 2(e) and 2(f)). With a smaller learning rate such as 1.0e-7, though the model’s performance on MMLU and IFEval can be well-preserved, the model’s performance on table tasks such as FEVEROUS is suboptimal under the same setup (66.86 compared to 74.63 at a learning rate of 5.0e-6). In contrast, when the learning rate is too large, such as 1.0e-5, we observe a significant decline in the model’s performance on both MMLU and IFEval, suggesting that a larger learning rate may hurt the model’s general capabilities. We note that all the existing table LLMs use a large learning rate of 2e-5 (Table 1), which explains their compromised out-of-domain table understanding ability and general capabilities compared to their base models in Table 3.
@@
We observe in Figure 2 that on table tasks such as FeTaQA and HiTab, there is a period where the model’s performance boosts up quickly, typically happening when tuning on the first 200 examples. Later, the performance improvement appears marginal. This aligns with the findings of Zhou et al. (2024) that the foundational performance of the LLM can be improved with a limited amount of high-quality data in the instruction tuning stage. We hypothesize that with the first few hundred examples, the model is able to enhance its table reasoning ability quickly. After this point, the model’s performance increase may primarily come from fitting the nuanced patterns in these datasets. Therefore, unlike the existing table LLMs that may involve up to two million training instances as seen in Table 1, we choose to train on 200 instances for each dataset in Section 4.
In addition, we can achieve competitive or even SOTA performance with limited data. On HiTab, with a learning rate of 1.0e-6 and 1,500 examples, we achieve an accuracy score of 66.29, outperforming the previous SOTA performance of 64.71 by TableLLaMA. On FEVEROUS, with 1,500 examples, we achieve a better score of 74.63 compared to 73.77 by TableLLaMA. Although the credit also comes from the LLaMA 3.1 Instruct model, which is much stronger compared to the LLaMA 2 model that TableLLaMA is tuned from, we highlight that TableLLaMA has used two million data points in its table instruction tuning stage, including the entire training set of TabFact, FeTaQA, and HiTab, while here we use around 7% of the entire training data for HiTab. Our analysis demonstrates that with a strong foundational model and a good choice of learning rate, we can achieve competitive performance on table understanding tasks with limited training instances.
Model Learning Rate
Llama 2 7B Instruct 1.0e-6 / 5.0e-7
Llama 3.1 8B Instruct 1.0e-6 / 5.0e-7
QWen 2.5 7B Instruct 1.0e-6 / 5.0e-7
Mistral v0.3 7B Instruct 5.0e-7 / 1.0e-7
Phi 3 small 8K Instruct (7B) 5.0e-6 / 1.0e-6
@@
Full Analysis.
We provide a pointer to our full analysis and a short summary of the findings here.
  1. 1.
    Learning rate across LLMs under full parameter setup (Section C.4), and LoRA and QLoRA setups (Section C.5).
    Finding: We list our recommended learning rate per model in Table 4 for the full parameter setup. Please refer to Section C.5 for recommendations for LoRA and QLoRA setups.
  2. 2.
    Number of examples across LLMs under full parameter tuning setup (Section C.4), and LoRA and QLoRA setups (Section C.5).
    Finding: Under all these setups, there is a diminishing return as the number of training examples increases.
  3. 3.
    Number of epochs (Section C.2).
    Finding: We do not see significant performance gains when we increase the number of epochs. Therefore, we choose to train our model for two epochs in Section 4.
  4. 4.
    Multi-task training (Section C.3).
    Finding: There are synergy effects on these table tasks, therefore, we decide to fine-tune our model on a diverse range of tasks in Section 4.
In addition, we provide analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7.
4 TAMA
Based on our findings from Section 3, we build our general table understanding model TAMA by instruction tuning the LLaMA 3.1 8B Instruct model.
4.1 Experimental Setup
@@ 4.3 Hindsight Analysis
As shown in Figure 3(a), in the table understanding tasks, the learning rate of 1.0e-6 and 5.0e-7 yields the best overall performance, which coincides with our findings in Section 3. In addition, the model achieves its best aggregated performance around two to three epochs for both learning rate.
@@
For MMLU, both 1.0e-6 and 5.0e-7 maintain their performance, sometimes even slightly better than the original LLaMA 3.1 8B Instruct model. As revealed in Table 9, the performance boost is more pronounced on STEM category. We hypothesize that this is because table-related tasks typically involve data analysis that requires math reasoning, which belongs to the STEM category. Therefore, training on table-related tasks would lead to better STEM performance. This also explains the performance boost for MMLUPro in Table 8.
@@ 5 Related Work
The past decade has witnessed a paradigm shift in approaches to table understanding. Before the advent of LLMs, researchers typically adapt model structures to better interpret table data (Lebret et al., 2016; Liu et al., 2018; Yang et al., 2022). As language models demonstrate promising performance on various tasks (Devlin et al., 2019), researchers gradually shift their attention towards data-driven methods for table understanding. For instance, Yin et al. (2020); Herzig et al. (2020) pre-train BERT (Devlin et al., 2019) or BERT-derived model on large-volume of table data from sources such as Wikipedia to acquire better table representations. Xie et al. (2022) reveal the synergy effects of various structured tasks, including many table tasks, laying foundations to build a generalist model for structured data. In the era of LLMs, as LLMs possess innate table-understanding abilities, researchers also explore prompt engineering techniques to optimize LLMs for table tasks (Chang & Fosler-Lussier, 2023; Deng et al., 2024).
@@
Recently, researchers have increasingly focused on instruction tuning to enhance LLMs’ table understanding ability. As demonstrated by Touvron et al. (2023); Dubey et al. (2024); Chung et al. (2024); Su et al. (2024), instruction-tuning can improve model performance and generalization to unseen tasks. Meanwhile, models from the open-source LLaMA family (Touvron et al., 2023) demonstrate strong capabilities, leading researchers to instruction-tune these models for better table understanding. For instance, TableLLaMA (Zhang et al., 2024a) is instruction-tuned from a variant of LLaMA 2 model (Touvron et al., 2023), TableLLM (Zhang et al., 2024b) is instruction-tuned from CodeLLaMA, Wu et al. (2024) instruction-tune various foundational models such as LLaMA 3.1 (Dubey et al., 2024), resulting in their TableBenchLLM model. Moreover, Zheng et al. (2024a) treat tables as images and instruction-tune Vicuna (Chiang et al., 2023), a vision model that is originally fine-tuned from the LLaMA model, for table understanding. However, as revealed by Zheng et al. (2024a); Deng et al. (2024), treating tables as texts rather than images yields better performance. In this paper, we focus on table instruction tuning with tables fed as texts.
@@ 6 Conclusion
In this paper, we highlighted the limited out-of-domain table understanding ability and limited general capabilities of existing table LLMs. From our analysis, we found that the commonly adopted hyperparameters in existing table LLMs are suboptimal, and hyperparameter choices in table instruction tuning are crucial in shaping the model’s capabilities. Based on our analysis, we selected a set of hyperparameters and fine-tuned our own model, TAMA. Notably, as an 8B model, TAMA demonstrates strong table understanding ability, outperforming GPT-3.5 on most of the table understanding benchmarks, even achieving performance on par or better than GPT-4. Moreover, TAMA preserves strong general capabilities. We hope our findings and our model TAMA can facilitate future research on structured data.
Limitations and Future Directions
Scope of the Study.
Due to the space constraint, we provide further analysis across LLMs in Section C.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section C.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7. However, due to the scope of one study, we cannot exhaust every possible foundational model and every possible dataset.
Better Hyperparameter Selection Methods.
While our paper effectively demonstrates the importance of hyperparameter selection, we primarily rely on manual hyperparameter tuning. We encourage future efforts on automated approaches to hyperparameter optimization, specifically tailored for balancing domain adaptation and general capabilities.
Better Data Selection Methods.
Our work employs random sampling from datasets, which, while proving effective, suggests redundancy in current tableQA datasets. We encourage future efforts to investigate more principled data selection strategies based on diversity, difficulty, or information content to further improve efficiency.
@@
Acknowledgements
We leverage the LLaMA Factory library for model training and inference (Zheng et al., 2024b). We thank the anonymous reviewers for their valuable feedback and suggestions. We thank Yulong Chen, Xin Liu, Jie Ruan, Jung-Chun Liu for their feedback on the initial draft.
This project was partially funded by a grant from OpenAI and a Microsoft Foundational Model grant. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of OpenAI or Microsoft.